OpenStudy (anonymous):

Rationalize the denominator. 2sqrt(a)/7sqrt(6b)

OpenStudy (anonymous):

\[2\sqrt{a}\div7\]

OpenStudy (anonymous):

\[7\sqrt{6b}\]

OpenStudy (anonymous):

\[\frac{ 2\sqrt{a} }{ 7\sqrt{6b} }\]

OpenStudy (anonymous):

?